40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2826     Entry time: Tue Nov 9 19:01:43 2021
Author: aaron 
Type: Update 
Category: Laser 
Subject: fiber components to fiber box 

I mounted the fiber components (north and south Rio lasers, Faraday isolators, 90-10 beamsplitters, 50-50 beamsplitter, fiber EOM, 1611FC-AC) inside our breadboard-in-a-box from Thorlabs (attachment 1). 

Along the way

  • Inspected and cleaned all fiber tips that I connected. Fibers that are not connected have not necessarily been cleaned. I took some photos using my phone and a 15x macro lens (attachment 2 is of a clean fiber), but the camera kept focusing on some schmutz on the microscope's eyepiece, so they're not very illuminating. 
  • Turned off the S laser that Shruti was using, disconnected its cables, and reconnected the cables through the DB9 feedthrough on the box
  • Removed from cryo cavs table the beamsplitter, photodiode, and fiber connectors that we were borrowing for the three corner hat measurement. These are now inside the box.
  • labelled all fiber ends
  • Used the 10m patch cable that was running from cryo cavs table to PSOMA table to carry the beam from the fiber box to the beam launch on the PSOMA table. Note that one end of the patch cable had dust or damage on its face over the fiber core that I couldn't manage to clean off. Could try again, or get another (shorter) patch cable). 

Though I set up the beamsplitters in the box approximately where we will eventually want them, the beam currently does not pass through any beamsplitters. I left the optical layout for the south laser identical to how it was on the table, with the addition of a patch cable between the EOM and the launch: laser -> faraday -> EOM -> patch cable -> launch. The north laser is not connected to its drivers. 

Attachment 1: IMG_2187.jpg  2.991 MB  | Hide | Hide all
Attachment 2: IMG_2184.jpg  766 kB  | Hide | Hide all
ELOG V3.1.3-